13) (4 points) Write a MATLAB program (Be sure to begin with a block of comments including name, course number, final exam, p11)(4 points) Find Ix and Iy using mesh equations. Ix = Iy 20 50 200 V Ix 40 + 70 ly
ANSWER PROBLEM 13 NOT 11, PROBLEM 11 IS FOR PROBLEM 13. PLEASE FOLLOW DIRECTIONS.
Answer #1

20 W OS w 200 о Tv Apply mesh analysis for loop in -200 + 2017 +40 (da - Iv) = Go In - 40 TV =20070 Apply mesh analysis for l4G 1:33 AM 69 5 Δ MATLAB Drive > Aa.m* 1 A=[60 -40; 2 -40 160]; 3 B= 200;0]; 4 X= inv(A); 5 I=X*B; 6 display (I)4G 1:33 AM 69 >> Aa I - 4.0000 1.0000 >>